# Benchmarks and applications of the nuclear deexcitation event generator NucDeEx

### v2

@@ Abstract
Neutron multiplicity is a key observable in recent neutrino experiments that can enhance the sensitivity of various neutrino physics searches. Nuclear deexcitation plays a significant role in neutron emissions associated with neutrino-nucleus interactions; therefore, precise prediction of this process is eessential. To address this need, a general-purpose nuclear deexcitation event generator NucDeEx was developed and released as an open-source package. The treatment of low-lying discrete excited states was updated to better reproduce experimental data. Benchmarks were conducted using existing nuclear deexcitation event generators and experimental data. Application to other simulators, neutrino event generator NEUT and general particle simulation tool Geant4, are also presented.
@@ I Introduction
Neutron multiplicity has emerged as a key observable in recent neutrino detectors. Large liquid scintillator detectors such as KamLAND [1] and JUNO [2] are sensitive to gamma rays from nuclear deexcitation and thermal neutron capture because of their high light yield. In water Cherenkov detectors such as Super-Kamiokande (SK), the detection efficiency of the 2.2 MeV gamma ray from thermal neutron capture on hydrogen is limited to about 20% due to its small light yield [3]. Super-Kamiokande Gadolinium (SK-Gd) [4, 5] successfully enhanced the detection efficiency for the thermal neutron capture by dissolving Gd in water, which emits gamma rays with a total energy of approximately 8 MeV. Precise prediction of neutron multiplicity associated with neutrino–nucleus interactions is particularly important in searches for the diffuse supernova neutrino background (DSNB) [6, 7, 8]. In their analyses, the signal is inverse beta decay ($\bar{\nu}_{e}+p\rightarrow e^{+}+n$) whereas the dominant background arises from atmospheric neutrino interactions.
@@
Nuclear deexcitation plays a significant role in neutron emissions associated with neutrino-nucleus interactions. KamLAND reported good agreement between the predicted and observed neutron multiplicities associated with atmospheric neutrino interactions [1]. Notably, they coupled their custom deexcitation event generator with the neutrino event generator NuWro [9] in this study. The T2K far detector (SK) analysis reported significant overprediction in neutron multiplicity of charged-current quasielastic (CCQE)-enriched sample [10]. However, their recent investigation on NCQE interactions observed that the overprediction is largely mitigated by using alternative secondary interaction models in detector simulations [11]. SK-Gd also published several results recently on neutron multiplicity measurements associated with atmospheric neutrino interactions, showing that the observed multiplicities fall within the wide range predicted by various models, particularly nuclear deexcitation models [12, 13]. The reason for overprediction of neutron multiplicity at SK(-Gd) has been gradually identified by these studies: the deexcitation model coupled with the Bertini cascade model [14], which had long been used in SK detector simulation, was overly simple and lacked sufficient descriptive power. Not limited to these experiments, SNO [15] and MINERvA [16, 17] measured neutron multiplicity, and extensive investigation on model variations in large liquid scintillator detectors was conducted mainly for JUNO [18, 19, 20]. Beyond neutrino interaction, nuclear deexcitation modeling is also essential for predicting nucleon decay in nuclei [21, 22, 23]. Since deexcitation emits not only neutrons but also gamma rays, it has a significant impact on the detector response in experiments such as SK [24].
@@
To maximize the performance of these neutrino detectors, a reliable and general-purpose nuclear deexcitation event generator that can be used in combination with neutrino event generators is essential. Several attempts have been made to incorporate nuclear deexcitation processes within the framework of neutrino event generators: for example, the use of ABLA [25, 26, 27, 28] within the INCL++ (the Liége intranuclear cascade model) [29, 30] in NuWro [31], and the PEANUT model in FLUKA [32]. However, no tool has yet become widely adopted in neutrino physics analyses. This highlights the need for a flexible and well-validated deexcitation event generator that can be readily integrated into simulations.
@@ II.1 Discrete excited states
Decays from the continuum excited states are described by the Hauser-Feshbach model [40], as implemented in the nuclear reaction code TALYS [41], while those from discrete excited states are generated according to the branching ratios decribed below.
@@
The gamma-ray branching ratios from the discrete excited states of $^{15}\text{N}^{*}$ and $^{15}\text{O}^{*}$ have also been updated (Tab. 1). For example, in the case of $^{15}\text{N}^{*}$, version 1 considered two excited states at 6.32 MeV and 9.93 MeV below the proton separation energy based on Ref. [21]. Experimental data, however, indicate a non-negligible contribution from the excited states at 5.27 MeV and 5.30 MeV [42, 43]. Thus, version 2 considers a 9% contribution from these excited states, treating them as a single level due to their small energy separation. The minor contributions from excited states at 8.31 MeV and 9.05 MeV, although observed in Ref. [42], are neglected here because of their minimal impact. The excited state at 10.70 MeV is treated as part of the continuum excited states in version 2. Due to the absence of experimental data, the branching ratios for 15O are inferred from those of 15N, assuming that states with the same spin-parity and similar excitation energy exhibit analogous branching ratios. The excited state at 9.61 MeV of $^{15}\text{O}^{*}$, which corresponds to the 9.93 MeV state of $^{15}\text{N}^{*}$, lies above the separation energy and is therefore not considered as a discrete excited state here. The branching ratios for $^{11}\text{B}^{*}$ and $^{11}\text{C}^{*}$ remain unchanged from Table I in Ref. [33].
@@ III Comparisons with other deexcitation event generators
For simulations using NucDeEx, INCL++/FB, INCL++/ABLAv3p, and G4PreCompoundModel, the excitation energy $E_{x}$ was calculated by subtracting the proton separation energy $S_{p}$ from the removal energy $\tilde{E}$ determined by the previous SF [38, 39]:
@@ III.1 Energy spectra predicted by generators
Particles emitted by the deexcitation process have relatively low energies of several MeV compared to those emitted by direct quasielastic (QE) interactions, which are typically around 100 MeV. For example, the typical excitation energies of $s_{1/2}$-hole states of $^{11}\text{B}^{*}$ and $^{15}\text{N}^{*}$ are approximately 20–40 MeV. In most cases, sequential decay occurs rather than the emission of a single high-energy particle. Therefore, most of the excitation energy is consumed by the separation energy, which typically about 10 MeV, leaving the emitted particles with several MeV. Fig. 2 shows neutron energy spectra predicted by deexcitation event generators. Energy spectra for other particles are shown in Figs. 710. Although it depends on the particles species and generators, it is observed that 40–50% of the decay particles have low energies below the detection threshold in the experiment at RCNP with a proton beam [55, 56, 42]. This highlights a limitation of normal kinematics experiments using a proton beam. On the other hand, inverse kinematics experiments using an ion beam, such as GSI [58] and SAMURAI [59], can measure decay particles without detection thresholds because the center-of-mass frame is boosted. Thus, inverse kinematics experiments are more effective for measuring the deexcitation process.
@@ III.2 Comparisons with experimental data
Figure 3 shows a comparison of relative branching ratios of $n$ and $d/\alpha$ in $^{12}\text{C}(p,2p)^{11}\text{B}^{*}$ reactions measured at GSI [58]. This experiment used 12C beam at $400$ MeV/u. Because the center-of-mass frame is boosted, even low-energy decay particles have sufficient energy in the laboratory frame to be detected without thresholds. The FB model tends to overestimate $\alpha$ particle emission, resulting in a inconsistent trend with the experiment and other generators: the relative branching ratio of deuterons and $\alpha$ particles exceeds that of neutrons.
@@
Figure 4 shows comparisons of absolute branching ratios measured in $^{12}\text{C}(p,2p)^{11}\text{B}^{*}$ and $^{16}\text{O}(p,2p)^{15}\text{N}^{*}$ reactions at RCNP [55, 56, 42]. This experiment used proton beam at 392 MeV. In this experiment, detection thresholds of 3.1–4.6 MeV were applied to the decay particle measurement, resulting in 40–50% inefficiency, as presented in Sec. III.1 and Appendix. The branching ratio for decay channel $i$, $Br_{i}$, is given by:
@@
To systematically evaluate the performace of these generators, chi-squared values are calculated for each generator (Table 3). The chi-squared is defined as
@@
where $\Delta Br_{i}^{Data}$ is the statistical uncertainty of the branching ratios for experimental data, $Br_{i}^{MC}$ and $Br_{i}^{Data}$ are the branching ratios for decay channel $i$ of MC and experimental data, respectively. Both single step and multistep decays are taken into account in this calculation. Note that only statistical uncertainty is considered since the systematic uncertainty of the experiment is not provided. Thus, only the relative values of chi-squared are meaningful for evaluating the performance of the generators. It is found that INCL++/ABLA exhibits an energy conservation problem in the decay channel of $^{11}\text{B}^{*}\rightarrow t+\alpha+\alpha$, leading to a sizable overestimation of the branching ratio for this channel. Although the branching ratio of tritons appears to be close to the experimental value, this agreement is attributed to the violation of energy conservation. Therefore, this is not considered to be indicative of physical accuracy. INCL++/FB overestimates the branching ratio for $\alpha$ particles, while underestimating neutron emissons. G4PreCompoundModel also overestimates the branching ratio for $\alpha$, mainly due to the FB model selected in low-excition energy region. As observed in Fig. 3, the FB model consistently overpredicts $\alpha$ emissions, which does not agree well with experimental data. Furthermore, G4PreCompoundModel underpredicts single-step proton decays, which also failes to reproduce the data. As a result, NucDeEx shows the best, or at least comparable, agreement with experiments in terms of $\chi^{2}$. Generators based on the HF model tend to reproduce the experimental data better than others. This is consistent with expectations since the HF model includes angular momentum conservation.
@@ IV.1 Geant4
E487 and E525 experiments [62, 63] measured gamma-ray energy spectra of $^{16}\text{O}(n,\gamma)$ inclusive reactions. These experiments were conducted at RCNP using neutron beams with energies of 30 MeV, 80 MeV, and 250 MeV on water target. The observed gamma-ray energy spectra were compared with various combinations of cascade and deexcitation models [61]. Among the standard cascade models provided by Geant4 (BERT, BIC, INCL++), INCL++ shows the best agreement with the experimental data. Replacing the default deexcitation model (G4PreCompoundModel) in INCL++ with NucDeEx led to further improvement in reproducibility across all neutron energies (Figs. 5 and 6). In particular, a significant improvement is observed at a neutron energy of 80 MeV. This energy region is important for detector simulations of atmospheric neutrino interactions, where the typical neutron energy is around 100 MeV [64]. NCQE interactions, which do not produce a charged lepton in the final state, can be a background in DSNB searches when the total energy of emitted gamma rays reaches approximately 20 MeV. This improvement at 80 MeV is expected to directly enhance background estimations in DSNB searches at SK, especially for those from NCQE intractions.
@@
Figure 6: The observed and predicted gamma-ray energy spectra from neutron-oxygen interaction at 80 MeV neutron beam energies. The black dots represent the experimental data of E487 [62], and the magenga dots represent the E487 data normalized with the E525 neutron flux at 250 MeV. The red and dark blue lines are Geant4 simulations using INCL++ coupled with G4PreCompoundModel and NucDeEx, respectively. The figures are from Ref. [61], and the chi-squared values are summarized in the paper.
@@ IV.2 NEUT
NucDeEx has been officially integrated into the neutrino event generator NEUT [35] and is available from version 5.9.0. Currently, it is limited to QE interactions and is intended for used with the SF model [47]. The NEUT-NucDeEx interface supports previous SF [38, 39] as well as the new carbon SF [44], which has recently been implemented into NEUT. The excitation energy is determined according to Eq (1), neglecting the effects of the final-state cascade interactions. In contrast, nuclear species are determined after the cascade by counting the outgoing nucleons. Furthermore, a relativistic distorted wave impulse approximation (RDWIA) model has recently been implemented into NEUT [65] In this model, the missing energy distribution for each shell is parameterized using a Gaussian function, closely resembling that of the SF. Therefore, NucDeEx is expected to be compatible with the RDWIA model as well.
@@
For example, it is expected to improve background estimation in DSNB searches. Ref. [18, 19, 20] discusses generator-dependent variations by combining a custom deexcitation model with existing neutrino event generators, NuWro and GENIE [66]. NEUT allows us to perform similar investigations using NucDeEx in its package. Another application would be in the context of missing energy measurement using neutrinos from kaon decay-at-rest reported by JSNS2 Collaboration [67]. It would be valuable to investigate how the predictions vary when NucDeEx is combined with the new carbon SF that has high missing energy resolution, which offers enforced connection of primary interaction and deexcitation.

### v3

@@ Abstract
Neutron multiplicity is a key observable in recent neutrino experiments that can enhance the sensitivity of various neutrino physics searches. Nuclear deexcitation plays a significant role in neutron emissions associated with neutrino-nucleus interactions. Therefore, precise prediction of this process is essential. To address this need, a general-purpose nuclear deexcitation event generator NucDeEx was developed and released as an open-source package. The treatment of low-lying discrete excited states was updated to better reproduce experimental data. Benchmarks were conducted using existing nuclear deexcitation event generators and experimental data. Application to other simulators, neutrino event generator NEUT and general particle simulation tool Geant4, are also presented.
@@ I Introduction
Neutron multiplicity has emerged as a key observable in recent neutrino detectors. Large liquid scintillator detectors such as KamLAND [1] and JUNO [2] are sensitive to gamma rays from nuclear deexcitation and thermal neutron capture because of their high light yield. In water Cherenkov detectors such as Super-Kamiokande (SK), the detection efficiency of the 2.2 MeV gamma ray from thermal neutron capture on hydrogen is limited to about 20% due to its small light yield [3]. Super-Kamiokande Gadolinium (SK-Gd) [4, 5] successfully enhanced the detection efficiency for the thermal neutron capture by dissolving Gd in water, which emits gamma rays with a total energy of approximately 8 MeV. Precise prediction of neutron multiplicity associated with neutrino-nucleus interactions is particularly important in searches for the diffuse supernova neutrino background (DSNB) [6, 7, 8]. In their analyses, the signal is inverse beta decay ($\bar{\nu}_{e}+p\rightarrow e^{+}+n$) whereas the dominant background arises from atmospheric neutrino interactions.
@@
Nuclear deexcitation plays a significant role in neutron emissions associated with neutrino-nucleus interactions. KamLAND reported good agreement between the predicted and observed neutron multiplicities associated with atmospheric neutrino interactions [1]. Notably, they coupled their custom deexcitation event generator with the neutrino event generator NuWro [9] in this study. The T2K far detector (SK) analysis reported significant overprediction in neutron multiplicity of charged-current quasielastic (CCQE)-enriched sample [10]. However, their recent investigation on NCQE interactions observed that the overprediction is largely mitigated by using alternative secondary interaction models in detector simulations [11]. SK-Gd also published several results recently on neutron multiplicity measurements associated with atmospheric neutrino interactions, showing that the observed multiplicities fall within the wide range predicted by various models, particularly nuclear deexcitation models [12, 13]. The reason for overprediction of neutron multiplicity at SK(-Gd) has been gradually identified by these studies. The deexcitation model coupled with the Bertini cascade model [14], which had long been used in SK detector simulation, was overly simple and lacked sufficient descriptive power. Not limited to these experiments, SNO [15] and MINERvA [16, 17] measured neutron multiplicity, and extensive investigation on model variations in large liquid scintillator detectors was conducted mainly for JUNO [18, 19, 20]. Beyond neutrino interaction, nuclear deexcitation modeling is also essential for predicting nucleon decay in nuclei [21, 22, 23]. Since deexcitation emits not only neutrons but also gamma rays, it has a significant impact on the detector response in experiments such as SK [24].
The reduction of systematic uncertainties from neutrino-nucleus interactions is becoming increasingly critical in next-generation experiments such as Hyper-Kamiokande [25] and DUNE [26]. These experiments require more accurate modeling of final-state particles and nuclear deexcitation products to reduce systematic uncertainties. Recent measurements of low-energy $\alpha$ and $\gamma$ particles in the ArgoNeuT and MicroBooNE liquid-argon time projection chambers [27, 28] indicate that such detectors can measure particles originating from nuclear deexcitation. Such measurements provide powerful benchmarks for models describing nuclear deexcitation.
@@
To maximize the performance of these neutrino detectors, a reliable and general-purpose nuclear deexcitation event generator that can be used in combination with neutrino event generators is essential. Several attempts have been made to incorporate nuclear deexcitation processes within the framework of neutrino event generators: for example, the use of ABLA [29, 30, 31, 32] within the INCL++ (the Liége intranuclear cascade model) [33, 34] in NuWro [35], the PEANUT model in FLUKA [36], and low-energy neutrino event generator MARLEY [37]. However, no tool has yet become widely adopted in neutrino physics analyses. This highlights the need for a flexible and well-validated deexcitation event generator that can be readily integrated into simulations.
@@ II.1 Discrete excited states
Decays from the continuum excited states are described by the Hauser-Feshbach model [45], as implemented in the nuclear reaction code TALYS [46], while those from discrete excited states are generated according to the branching ratios described below.
@@
The gamma-ray branching ratios from the discrete excited states of $^{15}\text{N}^{*}$ and $^{15}\text{O}^{*}$ have also been updated (Table 1). For example, in the case of $^{15}\text{N}^{*}$, version 1 considered two excited states at 6.32 MeV and 9.93 MeV below the proton separation energy based on Ref. [21]. Experimental data, however, indicate a non-negligible contribution from the excited states at 5.27 MeV and 5.30 MeV [47, 48]. Thus, version 2 considers a 9% contribution from these excited states, treating them as a single level due to their small energy separation. The minor contributions from excited states at 8.31 MeV and 9.05 MeV, although observed in Ref. [47], are neglected here because of their minimal impact. The excited state at 10.70 MeV is treated as part of the continuum excited states in version 2. Due to the absence of experimental data, the branching ratios for 15O are inferred from those of 15N, assuming that states with the same spin-parity and similar excitation energy exhibit analogous branching ratios. The excited state at 9.61 MeV of $^{15}\text{O}^{*}$, which corresponds to the 9.93 MeV state of $^{15}\text{N}^{*}$, lies above the separation energy and is therefore not considered as a discrete excited state here. The branching ratios for $^{11}\text{B}^{*}$ and $^{11}\text{C}^{*}$ remain unchanged from Table I in Ref. [38].
@@ III Comparisons with other deexcitation event generators
For simulations using NucDeEx, INCL++/FB, INCL++/ABLAv3p, and G4PreCompoundModel, the excitation energy $E_{x}$ was calculated by subtracting the proton separation energy $S_{p}$ from the removal energy $\tilde{E}$ determined by the previous SF [43, 44]
@@ III.1 Energy spectra predicted by generators
Particles emitted by the deexcitation process have relatively low energies of several MeV compared to those emitted by direct quasielastic (QE) interactions, which are typically around 100 MeV. For example, the typical excitation energies of $s_{1/2}$-hole states of $^{11}\text{B}^{*}$ and $^{15}\text{N}^{*}$ are approximately 20–40 MeV. In most cases, sequential decay occurs rather than the emission of a single high-energy particle. Therefore, most of the excitation energy is consumed by the separation energy, which is typically about 10 MeV, leaving the emitted particles with several MeV. Figure 2 shows neutron energy spectra predicted by deexcitation event generators. Energy spectra for other particles are shown in the Appendix. Although it depends on the particles species and generators, it is observed that 40%–50% of the decay particles have low energies below the detection threshold in the experiment with a proton beam at Osaka University’s Research Center for Nuclear Physics (RCNP) [60, 61, 47]. These experiments, which are referred to as normal kinematics experiments, therefore have a limitation in detecting low-energy decay particles. On the other hand, experiments using an ion beam, referred to as inverse kinematics experiments, such as GSI [63] and SAMURAI [64], can measure decay particles without detection thresholds because the center-of-mass frame is boosted. Thus, inverse kinematics experiments are more effective for measuring the deexcitation process.
@@ III.2 Comparisons with experimental data
Figure 3 shows a comparison of relative branching ratios of $n$ and $d/\alpha$ in $^{12}\text{C}(p,2p)^{11}\text{B}^{*}$ reactions measured at GSI [63]. This experiment used 12C beam at $400$ MeV/u. Because the center-of-mass frame is boosted, even low-energy decay particles have sufficient energy in the laboratory frame to be detected without thresholds. The FB model tends to overestimate $\alpha$ particle emission, resulting in an inconsistent trend with the experiment and other generators: the relative branching ratio of deuterons and $\alpha$ particles exceeds that of neutrons.
@@
Figure 4 shows comparisons of absolute branching ratios measured in $^{12}\text{C}(p,2p)^{11}\text{B}^{*}$ and $^{16}\text{O}(p,2p)^{15}\text{N}^{*}$ reactions at RCNP [60, 61, 47]. This experiment used a proton beam of 392 MeV. In this experiment, detection thresholds of 3.1–4.6 MeV were applied to the decay particle measurement, resulting in 40%–50% inefficiency, as presented in Sec. III.1 and the Appendix. The branching ratio for decay channel $i$, $Br_{i}$, is given by
@@
To systematically evaluate the performance of these generators, chi-squared values are calculated for each generator (Table 3). The chi-squared is defined as
@@
where $\Delta Br_{i}^{Data}$ is the statistical uncertainty of the branching ratios for experimental data, $Br_{i}^{MC}$ and $Br_{i}^{Data}$ are the branching ratios for decay channel $i$ of MC and experimental data, respectively. Both single step and multistep decays are taken into account in this calculation. Note that only statistical uncertainty is considered since the systematic uncertainty of the experiment is not provided. Thus, only the relative values of chi-squared are meaningful for evaluating the performance of the generators. It is found that INCL++/ABLA exhibits an energy conservation problem in the decay channel of $^{11}\text{B}^{*}\rightarrow t+\alpha+\alpha$, leading to a sizable overestimation of the branching ratio for this channel. Although the branching ratio of tritons appears to be close to the experimental value, this agreement is attributed to the violation of energy conservation. Therefore, this is not considered to be indicative of physical accuracy. INCL++/FB overestimates the branching ratio for $\alpha$ particles, while underestimating neutron emissons. G4PreCompoundModel also overestimates the branching ratio for $\alpha$, mainly due to the FB model selected in low-excitation energy region. As observed in Fig. 3, the FB model consistently overpredicts $\alpha$ emissions, which does not agree well with experimental data. Furthermore, G4PreCompoundModel underpredicts single-step proton decays, which also fails to reproduce the data. As a result, NucDeEx shows the best, or at least comparable, agreement with experiments in terms of $\chi^{2}$. Generators based on the HF model tend to reproduce the experimental data better than others. This is consistent with expectations since the HF model includes angular momentum conservation.
@@ IV.1 Geant4
E487 and E525 experiments [67, 68] measured gamma-ray energy spectra of $^{16}\text{O}(n,\gamma)$ inclusive reactions. These experiments were conducted at RCNP using neutron beams with energies of 30 MeV, 80 MeV, and 250 MeV on water target. The observed gamma-ray energy spectra were compared with various combinations of cascade and deexcitation models [66]. Among the standard cascade models provided by Geant4 (BERT, BIC, INCL++), INCL++ shows the best agreement with the experimental data. Replacing the default deexcitation model (G4PreCompoundModel) in INCL++ with NucDeEx led to further improvement in reproducibility across all neutron energies (Figs. 5 and 6). Note that Ref. [68] pointed out a potential issue in the measurement of proton beam current in E487. This may have resulted in an incorrect normalization of the neutron beam flux, as shown in Fig. 6. In particular, a significant improvement is observed at a neutron energy of 80 MeV. This energy region is important for detector simulations of atmospheric neutrino interactions, where the typical neutron energy is around 100 MeV [69]. NCQE interactions, which do not produce a charged lepton in the final state, can be a background in DSNB searches when the total energy of emitted gamma rays reaches approximately 20 MeV. This improvement at 80 MeV is expected to directly enhance background estimations in DSNB searches at SK, especially for those from NCQE interactions.
@@
Figure 6: The observed and predicted gamma-ray energy spectra from neutron-oxygen interaction at 80 MeV neutron beam energies. The black dots represent the experimental data of E487 [67], while Ref. [68] pointed out a potential issue in the flux normalization. The magenta dots, representing the E487 data normalized with the E525 neutron flux at 250 MeV, are also shown to avoid this issue. The red and dark blue lines are Geant4 simulations using INCL++ coupled with G4PreCompoundModel and NucDeEx, respectively. The figures are from Ref. [66], and the chi-squared values are summarized in the paper.
@@ IV.2 NEUT
NucDeEx has been officially integrated into the neutrino event generator NEUT [40] and is available from version 5.9.0. Currently, it is limited to QE interactions and is intended for used with the SF model [52]. The NEUT-NucDeEx interface supports previous SF [43, 44] as well as the new carbon SF [49], which has recently been implemented into NEUT. The excitation energy is determined according to Eq (1), neglecting the effects of the final-state cascade interactions. In contrast, nuclear species are determined after the cascade by counting the outgoing nucleons. Furthermore, a relativistic distorted wave impulse approximation (RDWIA) model has recently been implemented into NEUT [70] In this model, the missing energy distribution for each shell is parametrized using a Gaussian function, closely resembling that of the SF. Therefore, NucDeEx is expected to be compatible with the RDWIA model as well.
@@
For example, it is expected to improve background estimation in DSNB searches. Reference [18, 19, 20] discusses generator-dependent variations by combining a custom deexcitation model with existing neutrino event generators, NuWro and GENIE [71]. NEUT allows us to perform similar investigations using NucDeEx in its package. Another application would be in the context of missing energy measurement using neutrinos from kaon decay-at-rest reported by JSNS2 Collaboration [72]. It would be valuable to investigate how the predictions vary when NucDeEx is combined with the new carbon SF that has high missing energy resolution, which offers enforced connection of primary interaction and deexcitation.